OpenStudy (anonymous):

what is 2x7 - 10/9-4. please have a great answer because this is due tom. i'll give you a medal if you do

OpenStudy (campbell_st):

looks like order of operation 2 x 7 = 14 10/9 = 1 1/9 so you have 14 - 1 1/9 - 4 = 14 - 5 1/9 = 8 8/9